### v1

# Extending Graph Condensation to Multi-Label Datasets: A Benchmark Study

Abstract
As graph data grows increasingly complicate, training graph neural networks (GNNs) on large-scale datasets presents significant challenges, including computational resource constraints, data redundancy, and transmission inefficiencies. While existing graph condensation techniques have shown promise in addressing these issues, they are predominantly designed for single-label datasets, where each node is associated with a single class label. However, many real-world applications, such as social network analysis and bioinformatics, involve multi-label graph datasets, where one node can have various related labels. To deal with this problem, we extends traditional graph condensation approaches to accommodate multi-label datasets by introducing modifications to synthetic dataset initialization and condensing optimization. Through experiments on eight real-world multi-label graph datasets, we prove the effectiveness of our method. In experiment, the GCond framework, combined with K-Center initialization and binary cross-entropy loss (BCELoss), achieves best performance in general. This benchmark for multi-label graph condensation not only enhances the scalability and efficiency of GNNs for multi-label graph data, but also offering substantial benefits for diverse real-world applications.
1 Introduction
Graph-structured data are fundamental to many real-world applications, including social networks, academic citation networks, chemical molecules, protein-protein interaction networks, mapping services, and product recommendation systems Battaglia et al. (2018); Wu et al. (2020); Zhou et al. (2020); Wu et al. (2022). In these graph structures, nodes represent entities (e.g., users in social networks), while edges represent relationships between these entities (e.g., social connections between users)Zhang et al. (2023).
Graph Neural Networks (GNNs) are a class of deep learning models specifically designed to process and learn from graph-structured data. It utilizes the topology of the graph to capture dependencies and learn meaningful nodes, edges, and graph-level representations. By extending deep learning to non-Euclidean domains, GNNs have achieved state-of-the-arts performance in many graph machine learning applications Zhang et al. (2023); Hamilton et al. (2018). For example, in the recommendation system Li et al. (2019); Mao et al. (2021); Wu et al. (2022); Zhang et al. (2019a), drug discoveries Jiang et al. (2021); Duvenaud et al. (2015); Zhang et al. (2021), fraud account detection Dou et al. (2020); Wang et al. (2021) and traffic forecasting Jiang & Luo (2022). GNNs combine the node features and structure of graph as the input and output the representations of graph information Hamilton et al. (2018); Velickovic et al. (2017); Kipf & Welling (2016). Among all the GNNs techniques, the most broadly followed methods is recursive neighborhood aggregation scheme, which aims to get the representation of each node, respectively Xu et al. (2018). By applying different neighborhood aggregation and graph-level pooling methods, GNNs are capable of handling various of tasks like node classification Xiao et al. (2022), link prediction Hasan & Zaki (2011), and graph classification Zhang et al. (2018).
Even though GNNsZhang et al. (2023; 2019b) are widely used for analyzing graphs, training these models on large-scale graphs is still computationally expensive Gao et al. (2018); Bojchevski et al. (2020). Furthermore, nowadays, graph data is becoming gigantic with numerous nodes and edges in a single graph Zhang et al. (2023). For example, the Twitter user graph has 288M monthly active users as of 3/2015 and an estimated average of 208 followers per user for an estimated total of 60B followers (edges) Ching et al. (2015). Such industry graphs can be two orders of magnitude larger, like hundreds of billions or up to one trillion edges. In other words, the complexity and large scale of graph data pose significant challenges for training GNNsJin et al. (2022b). Despite the scalability challenges, for some specific applications that need re-training multiple times models would encounter more hardships rather than computation cost.
To address these large-scale datasets challenges, graph condensation methods recently Xu et al. (2024); Gao et al. (2024) have gained lots of attention. Graph condensation is motivated by condensing a large graph dataset into a smaller informative synthetic graph, thereby enhancing the scalability and efficiency of GNNs Xu et al. (2024); Gao et al. (2024); Hashemi et al. (2024). The goal is to achieve comparable performance in GNNs using this synthetic small graph, instead of relying on the original large-scale graph Loukas (2018); Jin et al. (2022a). By doing so, condensed synthetic graph enables GNNs to maintain strong predictive performance while significantly reducing the computational resources required for tasks such as training and inference. By eliminating redundant information in original dataset, it makes the synthetic graph more manageable within the constraints of limited computation resources, thereby providing better support for graph data mining tasks and applications such as Continual learning Liu et al. (2023), Network Architecture Search (NAS)Gao et al. (2021), etc. Moreover, take node classification task as an example, the node can be well classified is because GNNs have learned to capture the unique pattern of nodes to distinguish them from other nodes in different classes.
Existing graph condensation works Xu et al. (2024); Gao et al. (2024) investigate different condensing strategies including gradient distance matchingJin et al. (2022b); Yang et al. (2023), trajectory matchingZheng et al. (2024); Zhang et al. (2024), kernel ridge regressionXu et al. (2023); Wang et al. (2024), and distribution matchingLiu et al. (2022; 2023). Even though they have achieved great success, the original and synthetic datasets are all single-label ones. In many real-world scenarios, the nodes of graph can be associated with multiple labels in lots of situations Huang & Zhou (2012); Read et al. (2011). For instance, users are allowed to join multiple groups that respectively represent their diverse interests. Therefore, nodes of social networks Shi et al. (2019) are annotated with several tags profiling users’ preferences in different domains. Similarly, a paper in a citation network Akujuobi et al. (2019) may be associated with multiple research topics, and in the protein-protein interaction networks Zeng et al. (2019), it is used to classify protein functions in biology, which are associated with multiple biological processes, like molecular activities or cellular components. These functions are typically inferred based on how proteins interact with one another, with each interaction providing insight into the role of protein within the cell, such as catalyzing reactions, transporting molecules, or transmitting signals. So in the application of graph datasets, multi-label node classification is a fundamental and practical task in graph data miningShi et al. (2019; 2020); Akujuobi et al. (2019); Zhao et al. (2023). In the multi-label graph scenario, each node not only has content (or features) but also associated with multiple class labels. The goal is to assign one or more labels to each node. However, existing graph condensation techniques are designed to address the single-label graph scalability issue, typically condensing the large graph based on class-wise condensing. In this way, it can be hard for us to get the synthetic small graph from the multi-label graph datasets. While directly applying single-label condensation methods to multi-label graphs would result in a loss of critical information, as the class-wise assumption does not hold. This reminds as the open challenge.
To bridge this gap from single-label to multi-label setting, in this work, we adapt available graph condensation techniques to the mainstream multi-label datasets. Our approach includes modifications to the synthetic dataset initialization and original datasets condensation matching stages. By evaluating various adaptation settings, we identify three key insights into the optimal combination of initialization and optimization methods. Using these best settings, we further analyze graph condensation models and present three additional observations, highlighting the unique challenges of condensing multi-label graph datasets.
Our contributions are as follows:
  1. 1.
    We extend classic SOTA graph condensation methods, including GCond, GCDM, and SGDD, to the multi-label graph dataset scenario by introducing multi-label synthetic dataset initialization and condensing optimization methods.
  2. 2.
    To find the best adaption strategies, we compare different initialization techniques such as Random sampling, Herding, K-Center and probability synthetic multi-label combine with two different multi-label loss functions – SoftMarginLoss and BCELoss.
  3. 3.
    With the best adaption settings, we evaluate the condensation methods with F1-micro and F1-macro scores on eight real-world multi-label datasets: PPI, PPI-large, Yelp, DBLP, PCG, HumanGo, EukaryoteGo, and OGBN-Proteins. Finally, we find the GCond method generally works best with K-Center initialization and BCELoss.
2 Related Work
2.1 Graph Neural Networks
Graph Neural Networks (GNNs) are a class of deep learning models specifically designed to process and learn from graph-structured data. GNNs have achieved outstanding performance in many graph machine learning applications Zhang et al. (2023); Hamilton et al. (2018). For instance, in the recommendation system on social network Li et al. (2019); Mao et al. (2021); Wu et al. (2022); Zhang et al. (2019a), drug discoveries from molecule graphs Jiang et al. (2021); Duvenaud et al. (2015); Zhang et al. (2021), fraud account detection on financial graphs Dou et al. (2020); Wang et al. (2021) and traffic forecasting in transportation graph Jiang & Luo (2022).
GNNs combine the node features and structure of graph as the input and output the representations of graph Hamilton et al. (2018); Velickovic et al. (2017); Kipf & Welling (2016). Among all the GNNs techniques, they broadly follow a recursive neighborhood aggregation scheme to get the representation of each node Xu et al. (2018). Some of the typical GNNs models like the Graph Convolutional Network (GCN) Kipf & Welling (2016) operate by aggregating and transforming information from neighboring nodes layer by layer. Variants like Graph Attenstion Nesworks (GAT) Velickovic et al. (2017) use attention mechanisms to prioritize certain neighbors, while GraphSAGE Hamilton et al. (2017) samples neighborhoods to scale to large graphs. By applying different neighborhood aggregation and graph-level pooling methods, GNNs are capable of handling various of tasks Xiao et al. (2022); Hasan & Zaki (2011); Zhang et al. (2018) like node classification, link prediction, and graph classification.
2.2 Multi-Label Classification on Graphs
Multi-label classification on graphs is a fundamental task in graph machine learning, where nodes in a graph are associated with multiple labels simultaneously. This task is critical in numerous real-world applications, including social networks, bioinformatics, and recommendation systems. For instance, in social networks, users may belong to multiple interest groups, while in bioinformatics, proteins may participate in multiple biological processes such as molecular functions and cellular activities Huang & Zhou (2012); Read et al. (2011); Shi et al. (2019); Akujuobi et al. (2019). In multi-label classification, the goal is to predict a set of labels for each data point in the dataset. The labels can be thought of as binary variables, where a label is either present (1) or absent (0). Recently, there have been several approaches that use neural networks for multi-label classification.
Traditional multi-label classification methods relied on adapting single-label classification techniques to multi-label scenarios. Popular strategies include problem transformation methods (e.g., binary relevance and classifier chains) and algorithm adaptation approaches, which extend existing classifiers to handle multiple labels directly Read et al. (2011); Tsoumakas & Katakis (2008). These approaches use architectures such as recurrent neural networks (RNNs) and convolutional neural networks (CNNs) to learn representations of the input data and make predictions for the labels. However, these approaches were not designed to exploit the graph structure and are therefore suboptimal for graph data.
With the advent of GNNs, researchers began exploring multi-label classification directly on graph-structured data. GNNs excel in capturing the relational and topological information of graphs, making them well-suited for multi-label classification tasks. For example, GraphSAGE Hamilton et al. (2017) and GAT Velickovic et al. (2017) utilize neighborhood aggregation to generate node embeddings, which can be fed into multi-label classifiers. Specialized models like ML-GCN Chen et al. (2019) integrate graph convolutional layers with multi-label prediction heads to improve performance on multi-label datasets.
2.3 Graph Condensation
Despite the impressive performance of GNN models, training them on large-scale datasets is often challenging due to high computational costs, especially when retraining or dealing with redundant information. To address these issues, graph condensation techniques have been developed. These methods distill a large-scale graph into a smaller yet informative synthetic graph, improving the scalability and efficiency of GNNs Xu et al. (2024); Gao et al. (2024); Hashemi et al. (2024). The condensation process generates a compact representation that preserves essential topological and feature-based properties from the original graph Loukas (2018); Jin et al. (2022a). This allows GNNs to achieve comparable performance using the condensed graph while significantly reducing the computational resources needed for training and inference.
Existing graph condensation works Xu et al. (2024); Gao et al. (2024) investigate different condensing strategies including gradient distance matchingJin et al. (2022b); Yang et al. (2023), trajectory matchingZheng et al. (2024); Zhang et al. (2024), kernel ridge regressionXu et al. (2023); Wang et al. (2024), and distribution matchingLiu et al. (2022; 2023). Condensed graphs facilitate more efficient model execution without sacrificing the quality of learning, making this technique a promising method for real-world GNN deployments. By eliminating redundant information in original dataset, graph condensation makes the synthetic graph more manageable within the constraints of limited computation resources, thereby providing better support for graph data mining tasks and applications such as Continual learning Liu et al. (2023) and Network Architecture Search (NAS)Gao et al. (2021), etc. Moreover, take node classification as an example, the reason a node can be well classified is that GNNs have learned to capture the unique pattern of nodes to distinguish them from other nodes in different classes.
3 Problem Formulation
Figure 1: Workflow for multi-label graph condensation. It shows the process of condensing a large multi-label graph $\mathcal{G}=\{A,X,Y\}$ into a smaller synthetic graph $\mathcal{S}=\{A^{\prime},X^{\prime},Y^{\prime}\}$, where $Y$ and $Y^{\prime}$ represent the multi-label matrix. Various matching strategies, denoted by $\mathcal{M}(\cdot)$, are employed to ensure that key information in the original graph is captured. Our goal is to use the synthetic graph $\mathcal{S}$ to train a GNN that achieves comparable performance to the one trained on the original graph $\mathcal{G}$, thus reducing graph size while retaining performance.
Consider a graph $\mathcal{G}=\{A,X,Y\}$, where $N$ is the number of nodes, $A\in\mathbb{R}^{N\times N}$ is the corresponding adjacency matrix, $X=\{x_{1},x_{2},\cdots,x_{N}\}\in\mathbb{R}^{N\times d}$ is the $d$-dimensional node feature matrix. In multi-label scenario, each node is associated with multiple labels revealing characteristics or semantics of the node. In this case, the label matrix can be represented as $Y\in\{y_{1},y_{2},\cdots,y_{N}\}\in\{0,1\}^{N\times K}$, where $K$ is the number of classes. Following the definition of graph condensation, our goal is to learn a smaller synthetic graph denoted as $\mathcal{S}$, which can contain the ability to train downstream tasks’ GNNs and achieve competitive performance compared with the original large-scale complex graph $\mathcal{G}$. Similarly, the synthetic graph can be denoted as $\mathcal{S}=\{A^{\prime},X^{\prime},Y^{\prime}\}$ along with $A^{\prime}\in\mathbb{R}^{N^{\prime}\times N^{\prime}}$, and $X^{\prime}\in\mathbb{R}^{N^{\prime}\times d}$, $Y^{\prime}\in\{0,1\}^{N^{\prime}\times K}$. During this stage, our expectation is that the scale of the synthetic graph can be much smaller compared with the original graph, denotes as $N^{\prime}\ll N$.
In this work, we adapt the graph condensation techniques to a multi-label setting, trying to get the synthetic multi-label graph. First, we introduce the general framework of graph condensation in Figure 1. Generally, graph condensation methods are using different matching strategies to optimize the synthetic graph $\mathcal{S}$ similar to the original graph $\mathcal{G}$. Despite these various strategies, the key point is to keep the comparable performance of GNNs while training on small graph. Therefore, our objective can be formulated as follows Jin et al. (2022b):
$\mathcal{S^{*}}=\arg\min_{\mathcal{S}}\mathcal{M}(\theta_{\mathcal{S}}^{*},\theta_{\mathcal{G}}^{*})\quad\text{s.t.}\quad\theta^{*}=\arg\min_{\theta}\mathcal{L}({GNN}_{\theta}),$ (1)
where ${GNN}_{\theta}$ denotes a GNN parameterized with $\theta$. Specifically, $\theta_{\mathcal{S}}$ and $\theta_{\mathcal{G}}$ are the parameters trained on graph $\mathcal{S}$ and $\mathcal{G}$, respectively. $\mathcal{M}(\cdot)$ is the matching strategy used to match the model $\theta_{\mathcal{S}}^{*}$ trained synthetic graph $\mathcal{S}$ to the one $\theta_{\mathcal{G}}^{*}$ trained on the original graph $\mathcal{G}$. $\mathcal{L}(\cdot)$ is the loss function used to measure the difference between model predictions and ground truth. For most graph condensation methods, the loss function is cross-entropy loss.
As the specific initialization of model parameters $\theta$ can lead to the overfitting problem Jin et al. (2022b), one common solution Wang et al. (2018) is to generate the synthetic data following a distribution of random initializations of $P_{\theta_{0}}$. Furthermore, the reformulated optimization problem as follows:
$\mathcal{S^{*}}=\arg\min_{\mathcal{S}}\mathbb{E}_{\theta_{0}\sim P_{\theta_{0}}}[\mathcal{M}(\theta_{\mathcal{S}}^{*},\theta_{\mathcal{G}}^{*})$ (2)
[…]
where $\theta(\theta_{0})$ indicates $\theta$ is a function acting on $\theta_{0}$.
4 Graph Condensation Methods
Building on the global understanding of existing single-label graph condensation methods, we observe that while different graph condensation methods differ primarily in their matching strategies, the overall workflow remains largely consistent. Therefore, before adapting these methods to the multi-label scenario, we will illustrate the different matching strategies of condensation methods first.
4.1 GCond: Gradient Matching
By computing the gradient of $\mathcal{L}$ w.r.t. $\mathcal{S}$ and optimize $\mathcal{S}$ via gradient descent. For each training step, the $opt_{\theta}$ is the update rule, consider the one-step gradient descent for model parameters:
$\theta^{\mathcal{S}}_{t+1}\leftarrow\theta^{\mathcal{S}}_{t}-\eta\nabla_{\theta}\mathcal{L}({GNN}_{\theta^{\mathcal{S}}_{t}}(A^{\prime},X^{\prime}),Y^{\prime})$ (3)
$\theta^{\mathcal{G}}_{t+1}\leftarrow\theta^{\mathcal{G}}_{t}-\eta\nabla_{\theta}\mathcal{L}({GNN}_{\theta^{\mathcal{G}}_{t}}(A,X),Y)$
where $\theta^{\mathcal{S}}_{t}$ and $\theta^{\mathcal{G}}_{t}$ denote the model parameters trained on $\mathcal{S}$ and $\mathcal{G}$ at step $t$. $\eta$ is the learning rate for the gradient descent.
By matching the training gradient trajectory, define distance function $D(\cdot,\cdot)$, define the gradient matching strategy $\mathcal{M}_{GCond}$ as:
$\mathcal{M}_{GCond}=D(\nabla_{\theta}\mathcal{L}({GNN}_{\theta^{\mathcal{S}}_{t}}(A^{\prime},X^{\prime}),Y^{\prime}),$ (4)
$\nabla_{\theta}\mathcal{L}({GNN}_{\theta^{\mathcal{G}}_{t}}(A,X),Y))$
Assume $T$ is the number of steps of the whole training, the final optimized objective is:
$\mathcal{S^{*}}=\arg\min_{\mathcal{S}}\mathbb{E}_{\theta_{0}\sim P_{\theta_{0}}}[\sum_{t=0}^{T-1}D(\nabla_{\theta}\mathcal{L}({GNN}_{\theta^{\mathcal{S}}_{t}}(A^{\prime},X^{\prime}),Y^{\prime}),$ (5)
$\nabla_{\theta}\mathcal{L}({GNN}_{\theta^{\mathcal{G}}_{t}}(A,X),Y))].$
Direct joint learning can be highly challenging because the three variables $A^{\prime},X^{\prime}$, and $Y^{\prime}$ in synthetic graph are interdependent. One straightforward way is to treat $A^{\prime}$ and $X^{\prime}$ as free parameters. By modeling the synthetic graph structure as a function of the condensed node features, synthetic graph structure $A^{\prime}$ can be learned by:
$A^{\prime}=g_{\phi}(X^{\prime})$ (6)
where $\phi$ is the parameter of a multi-layer neural network. So the rewrite objective is:
$\mathcal{S^{*}}=\arg\min_{X^{\prime},\phi}\mathbb{E}_{\theta_{0}\sim P_{\theta_{0}}}[\sum_{t=0}^{T-1}D(\nabla_{\theta}\mathcal{L}({GNN}_{\theta^{\mathcal{S}}_{t}}(g_{\phi}(X^{\prime}),X^{\prime}),Y^{\prime}),$ (7)
$\nabla_{\theta}\mathcal{L}({GNN}_{\theta^{\mathcal{G}}_{t}}(A,X),Y))]$
4.2 SGDD: Structure-broadcasting
Instead of treating $A^{\prime}$ and $X^{\prime}$ in the synthetic graph $\mathcal{S}$ as free parameters, recent work Yang et al. (2023) points out that this way ignores the original structure $A$ in $\mathcal{G}$. To address this problem, they broadcast $A$ as supervision for the generation of $A^{\prime}$. By introducing graphon Gao & Caines (2019); Ruiz et al. (2020); Xia et al. (2023) to matching the different shapes between $A$ and $A^{\prime}$, SGDD use random noise $\mathcal{Z}(N^{\prime})\in\mathbb{R}^{N^{\prime}\times N^{\prime}}$ as input coordinates. Through the generative model to synthesize an adjacency matrix $A^{\prime}$ with $N^{\prime}$ nodes, the process can be formulated as:
$A^{\prime}=GEN(\mathcal{Z}(N^{\prime});\Phi)$ (8)
where $GEN(\cdot)$ is the generative model with parameter $\Phi$. Then the structure optimization is:
$\mathcal{M}_{structure}=Dis(A,GEN(\mathcal{Z}(N^{\prime});\Phi)),$ (9)
where $A$ is supervision and $Dis(\cdot)$ is a metric that measure the difference between $A$ and $A^{\prime}$. By adding the corresponding nodes information from $X^{\prime}$ and $Y^{\prime}$ to the inherent relation learning Pfeiffer III et al. (2014); Shalizi & Thomas (2011), the revised generation model is $GEN(\mathcal{Z}(N^{\prime})\oplus X^{\prime}\oplus Y^{\prime};\Phi)$, $\oplus$ denotes the concatenate operation. In general, SGDD concurrently optimizes the parameters $X^{\prime}$ and $A^{\prime}$. While the refinement of $X^{\prime}$ is achieved through a gradient matching strategy, whereas the $A^{\prime}$ is enhanced using the Laplacian energy distribution(LED) matching technique Tang et al. (2022); Das et al. (2016); Gutman & Zhou (2006). During each step, the other component is frozen to ensure effective refinement, and the overall training loss function can be summarized as:
$\mathcal{M}_{SGDD}=\mathcal{M}_{GCond}+\alpha\mathcal{M}_{structure}+\beta{||A||}_{2},$ (10)
where ${||A||}_{2}$ is proposed as a sparsity regularization term, $\alpha$ and $\beta$ are trade-off parameters.
4.3 GCDM: Distribution Matching
Inspired by the distribution matching strategy in image dataset condensation Zhao & Bilen (2023), GCDMLiu et al. (2022) adapt the receptive field distribution matching method in GNN. In the graph dataset, with the reproducinf kernal Hilbert space $\mathcal{H}$ defined by GNNs and with in each class $c$, GCDM optmize the maximum mean discrepancy (MMD). Let $\Phi_{\theta}$ be an $L-$layer GNN model parameterized by $\theta$, the loss of GCDM is:
$\mathcal{M}_{GCDM}=\min_{\phi,X^{\prime}}\sum_{c=0}^{C-1}r_{c}\cdot\max_{\theta_{c}}\left\|\frac{1}{V_{c}}\sum_{i\in V_{c}}\text{emb}_{i}^{c}-\frac{1}{V_{c}^{\prime}}\sum_{j\in V_{c}^{\prime}}\text{emb}_{j}^{c}\right\|^{2}_{2},$ (11)
where the node embeddings are:
$\{\text{emb}_{i}^{c}\}_{i=0}^{N-1}\leftarrow\Phi_{\theta_{c}}(A,X)$ (12)
$\{\text{emb}_{j}^{c}\}_{j=0}^{N^{\prime}-1}\leftarrow\Phi_{\theta_{c}}(A^{\prime}=g_{\phi}(X^{\prime}),X^{\prime})$ (13)
$V_{c}$ and $V_{c}^{\prime}$ represent the node sets belong to the class $c$ of original graph $\mathcal{G}$ and synthetic graph $\mathcal{S}$, respectively. $r_{c}$ denotes the condensed class ratio for each class of the synthetic graph.
5 Datasets and Experiment Settings
In this section, we first present the multi-label graph datasets used in our benchmark investigation, accompanied by the detailed introduction of label information in real-world application. Next, we detail the evaluation metrics used to assess the effectiveness of the adaptation methods, providing a robust framework for performance comparison.
Datasets. Specifically, we employ eight real-world datasets including PPI, PPI-large, Yelp, DBLP, OGBN-Proteins, PCG, HumanGo, EukaryoteGo. The details of those datasets are introduced as follows:
  • PPI Zeng et al. (2019) is the Protein-Protein Interaction network. It classifies protein functions based on the interactions of human tissue proteins. Positional gene sets are used, motif gene sets and immunological signatures as features and gene ontology sets as labels (121 in total), collected from the Molecular Signatures Database Liberzon et al. (2015). And PPI-large is the larger version of PPI dataset.
  • Yelp Zeng et al. (2019) is processed from a public open dataset from yelp.com. The nodes represent active users and the edges are the relationship between the users. The multi-label of each node represents the types of business like Coffee & Tea, Flowers & Gifts, Tours, and so on. Node features contain the information of all the reviews and rates by users, generated by pre-trained Word2Vec model Church (2017).
  • DBLP Akujuobi et al. (2019) is the citation network extracted from DBLP. The nodes represent authors and edges are the co-authorship between the authors. Multi-label here indicates the four research areas like database, data mining, information and retrieval and artificial intelligence.
  • OGBN-Proteins from Open Graph Benchmark Hu et al. (2020) is an undirected, weighted, and typed (according to species) graph. Nodes represent proteins, and edges indicate different types of biologically meaningful associations between proteins, e.g., physical interactions, co-expression or homology. The task is to predict the presence of protein functions in a multi-label binary classification setup, where there are 112 kinds of protein functions denote as labels to predict in total. Here the protein functions refer to the specific biological activities or roles that a protein performs within a cell or organism. These functions are typically described using Gene Ontology (GO) database Huntley et al. (2015) or other biological annotation systems and can fall into different categories.
  • PCG Zhao et al. (2023) is Protein-Phenotype graph dataset, which focused on predicting phenotypes associated with proteins. Phenotypes refer to observable traits or characteristics of diseases, and identifying these associations can be valuable for clinical diagnostics or discovering potential drug targets. The dataset is constructed using protein-phenotype associations from the DisGeNET database Piñero et al. (2020), where each protein is linked to one or more phenotypes. The phenotypes are grouped into disease categories based on the MESH ontology Bhattacharya et al. (2011), and labels with fewer than 100 associated proteins are removed. The PCG includes 3,233 proteins as nodes and 37,351 protein-protein interactions as edges. The task is to predict multiple phenotype labels for each protein based on its interactions and features.
  • HumanGo Chou & Shen (2007); Liberzon et al. (2015) contains 3,106 proteins, each potentially associated with one or more of 14 subcellular locations as multi-label. The proteins are represented as nodes in a graph, with features generated from their sequences using pre-trained model Yang et al. (2020). Protein-protein interactions form the edges of the graph, with each edge representing the confidence of different interaction types. Different location associations come from the GO database Huntley et al. (2015), which assigns standardized terms to proteins based on their roles in biological processes (e.g., cell division, metabolism), molecular functions (e.g., enzyme activity, receptor binding), and cellular components (e.g., nucleus, membrane).
  • EukaryoteGo Chou & Shen (2007); Liberzon et al. (2015) follows a similar structure but focuses on eukaryote proteins. It includes 7,766 proteins and their interactions, with proteins being assigned to one or more of 22 subcellular locations. Like HumanGo, it uses protein sequences for node features and protein-protein interactions for graph structure.
We follow the predefined data splits from Zeng et al. (2019); Zhao & Bilen (2023); Hu et al. (2020); Chou & Shen (2007). 1 presents an overview of the datasets’ characteristics.
Table 1: Dataset Statistics
Dataset #Nodes #Edges #Avg.Edges #Features #Labels Train/Val/Test
PPI 14,755 225,270 15.27 50 121 0.66/0.12/0.22
PPI-large 56,944 818,716 14.38 50 121 0.79/0.11/0.10
Yelp 716,847 6,977,410 9.73 300 100 0.75/0.10/0.15
DBLP 28,702 68,335 2.38 300 4 0.60/0.20/0.20
OGBN-Proteins 132,000 39,000,000 295.45 8 112 0.66/0.16/0.18
PCG 3,233 37,351 11.55 32 15 0.60/0.20/0.20
HumanGo 3,106 18,496 5.96 32 14 0.60/0.00/0.40
EukaryoteGo 7,766 13,818 1.78 32 22 0.60/0.00/0.40
Metrics. To evaluate the performance of adapted graph condensation methods, we utilize two widely recognized metrics in multi-label classification: F1-micro and F1-macro.
  • F1-micro score aggregates the contributions of all classes to compute the average F1 score. It does this by first calculating the total true positives, false positives, and false negatives across all labels, and then using these aggregated values to compute a global precision and recall. The F1-micro score is particularly useful when dealing with imbalanced datasets, as it gives equal weight to each instance rather than each class. It is defined as:
[…]
    where:
    Precision $=\frac{\text{TP}}{\text{TP}+\text{FP}}$
    Recall $=\frac{\text{TP}}{\text{TP}+\text{FN}}$
  • F1-macro score computes the F1 score for each class independently and then averages these scores. This metric treats all classes equally, regardless of their frequency in the dataset. It is particularly useful for understanding the model’s performance across all classes, especially in scenarios where some classes may be underrepresented. The F1-macro score is defined as:
[…]
After introduce the single-label graph condensation methods, we will focus on adjusting key components to fit the multi-label task while preserving the core matching mechanisms. This ensures that our framework can be extended to other existing graph condensation techniques, allowing for flexible and efficient adaptation across different methods. This benchmark presents a thorough evaluation of the various settings and methods used for graph condensation in the multi-label scenario. We structure the benchmark results into three key parts: (1) identifying the best initialization and loss function settings, (2) comparing the full results of all graph condensation methods across multiple datasets, and (3) analyzing the results through multi-label class distribution and label correlation.
6 Adapt Existing Graph Condensation Methods for Multi-label Datasets
Datasets C-rate Random Herding K-Center Probability Whole Dataset
Subgraph Nodes Subgraph Nodes Subgraph Nodes
PPI 1.00% 47.05 40.49 40.93 39.47 47.77 36.60 40.04 51.26
Yelp 0.02% 34.21 30.35 32.77 30.96 35.60 31.46 20.95 37.97
DBLP 0.80% 63.56 56.03 41.54 41.02 60.36 55.62 41.35 87.55
OGBN-Proteins 0.1% 14.12 15.44 11.62 15.36 29.95 24.81 15.53 18.86
PCG 4% 13.94 19.01 27.76 16.35 25.13 22.18 21.69 42.26
Table 2: F1-Micro Score (%) of Coreset Method with Different Initialization Strategies
Datasets C-rate Random Herding K-Center Probability Whole Dataset
Without $A^{\prime}$ With $A^{\prime}$ Without $A^{\prime}$ With $A^{\prime}$ Without $A^{\prime}$ With $A^{\prime}$ Without $A^{\prime}$ With $A^{\prime}$
PPI 1.00% 48.55 50.17 46.71 47.80 35.99 48.38 44.62 43.31 51.26
Yelp 0.02% 30.16 29.53 30.48 32.23 34.50 34.90 21.33 22.10 37.97
DBLP 1% 52.10 57.48 45.95 47.12 52.56 60.77 48.24 48.56 87.55
OGBN-Proteins 0.10% 22.03 28.69 20.54 25.19 28.09 29.04 22.10 27.01 30.59
PCG 4% 20.50 23.79 22.14 25.38 19.88 22.36 22.07 27.21 42.26
Table 3: F1-Micro Score (%) of GCond Method with Random/Herding/K-Center/Probability Distribution Initialization with/without Learning from Structure for SoftMarginLoss.
Datasets C-rate Random Herding K-Center Probability Whole Dataset
Without $A^{\prime}$ With $A^{\prime}$ Without $A^{\prime}$ With $A^{\prime}$ Without $A^{\prime}$ With $A^{\prime}$ Without $A^{\prime}$ With $A^{\prime}$
PPI 1.00% 49.95 51.35 47.47 48.23 36.79 49.27 45.65 43.91 51.26
Yelp 0.02% 29.24 30.76 31.76 32.41 32.36 33.94 22.96 25.62 37.97
DBLP 1% 53.41 59.40 46.87 48.50 54.21 70.74 49.73 51.33 87.55
OGBN-Proteins 0.10% 24.65 28.88 21.45 26.21 28.15 29.40 21.73 27.11 30.59
PCG 4% 22.26 25.45 24.66 26.76 20.17 25.58 23.47 28.28 42.26
Table 4: F1-Micro Score (%) of GCond Method with Random/Herding/K-Center/Probability Initialization with/without Learning from Structure for BCELoss.
After the introduction of different graph condensation methods based on the matching strategies in Section 4 including GCond, SGDD and GCDM, we apply the above methods in multi-label graph datasets in this section. However, these methods could not directly be applied to condense multi-label graphs majorly due to two reasons: (1) The synthetic graph initialization is designed for single-class augmentation with pre-defined condensed labels. (2) The condensation methods depend on GNNs to match the original graph $\mathcal{G}$ and synthetic graph $\mathcal{S}$, the classification objectives are designed for single-label node classification. Therefore, in this section, we would like to adapt the existing graph condensation methods by investigating different adaptation strategies to address these two issues. In particular, we first describe the existing issues and the corresponding resolvents in Section 6.1. Then, we conduct experiments to investigate which strategies work best for the multi-label graph condensation tasks in Section 6.2.
6.1 Multi-label Adaptation
During condensation stage, among all the condensation methods, class-wise batch sampling is widely used for synthetic graph optimization. For instance, in GCond, they initialize the synthetic graph labels $Y^{\prime}$ by predefine the labels for each class with condensation rate $r$. While optimization, to further reduce the memory usage and simplify the process, they calculate the gradient matching loss for nodes from different classes separately. As matching the gradients from a single class is easier than from all the classes, most of condensation researches follow this class-wise sampling method. However, in contrast to a single-labeled graph, a multi-labeled graph has multiple class labels assigned to each node, so the class-wise initialization and optimization process need to be revised to the multi-label scenario. Furthermore, when adapting single-label graph condensation techniques to multi-label node classification, the choice of the loss function is crucial to ensuring that the model effectively learns the complex relationships inherent in multi-label tasks. In single-label graph condensation, methods typically optimize for cross-entropy loss, which is suitable for tasks where each node belongs to a single class. However, in multi-label classification, nodes can belong to multiple classes simultaneously.
Therefore, to better determine the best setting for multi-label graph condensation we investigate different initialization and optimization methods to evaluate their impact on preserving multi-label relationships and classification accuracy.
Our objective is to assess how these initialization methods and loss functions influence the performance of the condensed graphs in multi-label classification tasks. To sum up, we focus on two key areas of adaptation: initialization and optimization of the synthetic graph. These modifications allow us to capture and preserve the essential multi-label features of the original graph.
Issue 1: Initialization of Synthetic Multi-Label Graph. In single-label settings, the condensed label $Y_{single}^{\prime}$ is predefined using ground-truth class distribution and retrieves $X^{\prime}$ randomly from each class.
As most of the graph condensation methods adapt Eq.6 to learn synthetic graph structure $A^{\prime}$ from node features $X^{\prime}$, one straightforward method for multi-label adaptation is sampling the nodes from original datasets. Inspired by the heuristic sampling algorithms of baseline Coreset methods used in GCond, the sampled subgraph by different strategies can be used for the synthetic multi-label initialization to get a better performance.
As some of the coreset methods gain comparable performance as baseline Welling (2009); Farahani & Hekmatfar (2009); Sener & Savarese (2017), we apply those heuristic algorithms to extract $N^{\prime}$ nodes from the full set of nodes to obtain subgraphs $S^{\prime}$ with corresponding condensation rates as the initialization methods.
  • Random selection is sampling the nodes randomly, so the probability of selecting node $i$ is $P(i)=\frac{1}{N}$.
  • Herding Welling (2009) method is more structured compared with random selection strategy. It is based on iterative minimization of the distance between the mean of the selected nodes’ feature vectors and the mean of the entire graph feature vectors. Let $\mu_{\mathcal{G}}$ present the mean feature vector of the original graph:
    $\mu_{\mathcal{G}}=\frac{1}{N}\sum_{i=0}^{N-1}x_{i}\quad\text{and}\quad\mu_{\mathcal{S}}=\frac{1}{N^{\prime}}\sum_{j=0}^{N^{\prime}-1}x^{\prime}_{j}$ (17)
  • K-Center Farahani & Hekmatfar (2009); Sener & Savarese (2017) method aims to maximize the coverage of the original graph feature space by minimizing the maximum distance between unselected nodes and their closest selected node. Let $\mathcal{S}=\{s_{1},s_{2},\cdots,s_{N^{\prime}}\}$ be the selected node set. The goal is to select nodes such that for every unselected node $i$, the distance to the nearest selected node $s_{j}\in\mathcal{S}$ is minimized:
    $\min_{s_{1},\cdots,s_{N^{\prime}}}\max_{i\in\{1,\cdots,N\}}\min_{j\in\{1,\cdots,N^{\prime}\}}\left\|x_{i}-x_{s_{j}}\right\|_{2}$ (18)
    The selection process would start from an arbitrary node as the first center $s_{1}$ and then iteratively select the next node $s_{t}$ which is farthest away from the selected nodes. The process continues until $N^{\prime}$ nodes are selected.
  • In the multi-label scenario, each node has multiple labels, and it is also important to ensure that the selected subgraph maintains the label distribution of the original graph. Here, we introduce probabilistic synthetic multi-label method, for each class $k$, define the label distribution as the fraction of nodes in the graph that belongs to class $k$:
    $p_{k}=\frac{1}{N}\sum_{i=1}^{N}Y_{i,k}$ (19)
    The goal is to ensure that the synthetic graph $\mathcal{S}$ maintains a similar label distribution. So for each synthetic label $y_{j}^{\prime}$ in $Y^{\prime}\in{0,1}^{N^{\prime}\times K}$, the label of class $k$ has a $p_{k}$ probability to be 1. Then we choose cosine similarity matching for the node features based on the real labels from $\mathcal{G}$ and probability synthetic labels.
Above all, in multi-label settings, we adapt subgraph initialization and probabilistic synthetic multi-label methods for multi-label graph condensation. Besides, in graph condensation, the structure $A^{\prime}$ of synthetic graph is optimized from nodes features $X^{\prime}$. Based on the researches Ding et al. (2022); Zhu et al. (2021); Jin et al. (2022b), this "Graphless" learning process can be highly useful for downstream data analysis, which also shows competitive performance. Therefore, by comparing different initialization methods with or without learning the graph structure, we get the best settings for the adaptation.
Issue 2: Optimization about Multi-Label Classification Loss.
In single-label settings, while the condensed graph is matched in various ways, the fixed label distribution can be used directly for synthetic graph initialization. With predefined synthetic label $Y^{\prime}_{single}\in\{0,\cdots,C-1\}^{N^{\prime}}$, the synthetic node feature is randomly retrieved from the particular class. As calculate the matching loss for nodes from different classes separately,is easier than that from all classes. In other words, this single-label class-wise sample method will directly sample each class to optimize. Specifically, for a given class $c$, it will sample a batch of nodes of class $c$ from original graph $\mathcal{G}$ with a portion of neighbors, denotes as $(A_{c},X_{c},Y_{c})\sim\mathcal{G}$. For synthetic graph $S$, only sample the specific nodes of class $c$ without the neighbors. By doing so, the synthetic graph will use all the other nodes for aggregation stages during condensation, similarly, this process can be wrote as $(A_{c}^{\prime},X_{c}^{\prime},Y_{c}^{\prime})\sim\mathcal{S}$. Further in terms of details, for each class $c$, there is a subset of nodes $N_{c}$ that belong to this class. The loss is computed class-wise, for example, the cross-entropy loss for single node $j$ with ground truth label $y_{j}$ and predicted logit $z_{j,c}$ for class $c$ is:
$\ell_{CE}(z_{j},y_{j})=-\log(\frac{e^{z_{j,y_{j}}}}{\sum_{c=0}^{C-1}e^{z_{j,c}}})$ (20)
The single-label final loss would be formulated as:
$\mathcal{L}_{single-label}=\sum_{c=0}^{C-1}\sum_{j\in N_{c}}\ell_{CE}(z_{j},y_{j})$ (21)
However, in multi-label classification $Y^{\prime}\in\{y_{1}^{\prime},\cdots,y_{N}^{\prime}\}\in\{0,1\}^{N^{\prime}\times K}$, each node can associate with multiple classes simultaneously. Typically, the multi-label classification task is modeled as $K$ independent binary classification tasks, for each task $k$, the model predicts a logit $z_{j,k}$ and ground-truth label $y_{j,k}$ indicates whether node $j$ belongs to class $k$. In multi-label classification, SoftMarginLoss Cao et al. (2019) incorporates margin-based optimization, which is beneficial when dealing with imbalanced datasets and overlapping classes. In multi-label settings, certain labels may frequently co-occur or exhibit correlations. SoftMarginLoss introduces a margin between positive and negative label predictions, encouraging the model to differentiate between these labels more confidently.
$\mathcal{L}_{Softmargin}(z_{j,k},y_{j,k})=log(1+e^{-z_{j,k}\cdot y_{j,k}})$ (22)
This formulation is particularly useful in ensuring that the model not only predicts correct labels but does so with confidence, reducing the chances of ambiguous predictions near the decision boundary. On the other hand, Binary Cross-Entropy Loss Durand et al. (2019) is a widely-used objective function for multi-label classification, as it independently evaluates each label for each node in the graph. Then the task-wise loss for each binary classification task is computed using Binary Cross-Entropy (BCE) Durand et al. (2019) loss:
$\mathcal{L}_{\text{BCE}}(z_{j,k},y_{j,k})=-(y_{j,k}\log(\sigma(z_{j,k}))+(1-y_{j,k})\log(1-\sigma(z_{j,k}))),$ (23)
where $\sigma(z_{j,k})=\frac{1}{1+e^{-z_{j,k}}}$ is the sigmoid function.
6.2 Experiments and Results Analysis
In this section, after the investigation of different initialization and optimization strategies. We combine them together to find out the best settings for multi-label graph condensation adaptation. Additional results in terms of other metrics are presented in Appendix A.
In this section, we conduct experiments on GCond to investigate the effectiveness of different initialization and optimization strategies on the representative datasets and corresponding condensation rates. We first introduce the experiment settings with specific methods and datasets. In graph condensation, the condensation ratio (often denoted as C-rate) is a measure of how much the original graph has been reduced in size for training purposes. This ratio represents the percentage or fraction of nodes retained from the full dataset in the condensed graph. For example, if the C-rate is set to 1%, it means that the scale of synthetic graph $\mathcal{S}$ is only 1% of original graph $\mathcal{G}$. Then we statistically analyzed the experimental results and presented the comparison of different settings for multi-label graph condensation.
Settings. For dataset selection, we pick the most representative datasets to simplify the choosing stage, for instance, PPI-large is the larger version of PPI dataset, so we only test the settings in PPI. With the representative datasets, we choose the C-rate based on the number of nodes in synthetic graph to compare the performance. Take the PPI dataset as the example, C-rate equals 1.00% means the synthetic graph has 147 nodes compared with original 14755 nodes. For all the synthetic graphs, the nodes number is around 150.
Specifically, we report the performance of various initialization strategies include subgraph sampling( Random, Herding, K-Center) Welling (2009); Farahani & Hekmatfar (2009); Sener & Savarese (2017), and probabilistic label sampling for synthetic $Y^{\prime}$. After the initialization, the synthetic graph $\mathcal{S}$ will be optimized by gradient matching strategy in GCond. Then we compare different optimization loss functions, including SoftMarginLoss Cao et al. (2019) and BCE loss Durand et al. (2019), across diverse datasets with pre-choosen condensation rates. Through these experiments, we aim to identify the optimal combination of initialization and loss function that achieves the best trade-off between graph size reduction and model performance in multi-label tasks.
Results and Analysis. Finally, by comparing different settings of multi-label scenario on GCond framework, we present the corresponding results of above initialization and optimization with or without structure learning. Specifically, Table 2 shows the F1-micro scores for four initialization strategies (Random, Herding, K-Center, and Probability) applied to both subgraph and node-only settings. This comparison across datasets allows us to identify which strategies best capture label distributions in multi-label tasks. Then Table 3, 4 compare performance under SoftMarginLoss and BCELoss optimizations, respectively. Besides, for each initialization method we investigate the performance of synthetic graph with or without structure information. These results help to pinpoint effective combinations of initialization and loss functions.
After analysis of the overall performance with key observations, we report the best setting for multi-label adaptation to better generalization in other methods. By observing the experimental results, we draw three corresponding observation conclusions:
Observation 1: K-Center sampling achieves optimal initialization.
In the GCond framework, K-Center sampling consistently provides the highest F1-micro scores for dense, large-scale datasets such as PPI, Yelp, and OGBN-Proteins in Tables 2, 3, 4.
In contrast, simpler strategies like Random and Herding yield better results in datasets with lower density, such as DBLP and PCG.
Following initialization, we tested SoftMarginLoss and BCELoss for optimization. Table 3 and Table 4 are the comparison of different initialization methods optimized by SoftMarginLoss and BCELoss, respectively. Whole dataset indicates the performance of GNN trained by original graph $\mathcal{G}$. (a) and (b) are the comparison of learning condensed graph with or without structure $A^{\prime}$. We present the combined bar and line chart to show a more intuitive comparison of the differences between various combinations of initialization and optimization methods. Interestingly, we find the best performance distribution is consistence even using different loss functions. K-Center initialization with or without graph structure optimization gains the best performance using BCELoss. In particular, datasets like PPI and DBLP exhibit notable performance gaps that highlight the advantage of structure learning (per Eq.6) in enhancing multi-label generalization.
Observation 2: Preserving graph structure could improves performance of condensed large datasets. We examined the effects of learning graph structure versus using nodes alone for label initialization. Results show substantial performance improvements when the graph structure is preserved, especially in large-scale datasets While all the experiments are conducted with the comparison of learning synthetic graph with or without structure information, we find out that with the additional $A^{\prime}$ learning, the condensed graph would perform better in the large-scale datasets.
Before condensation the structure information already shows its effectiveness. For example, in the comparison of initialization methods present in Table 2 PPI with a sampling of the C rate of 1. 00% and the K-Center, retaining the structure achieves an F1 micro score of 47. 77%, compared to 36. 60% with nodes alone. During condensation process, optimization with the structure information learning, same pattern is observed across various methods (see Table 3, 4 for a comparison with or without $A^{\prime}$ learning process), demonstrating the critical role of structural information.
Observation 3: BCELoss provides better overall results When comparing loss functions, BCELoss generally yields higher F1-micro scores, especially when combined with structure learning. For example, In Table 4, the DBLP dataset with K-Center sampling and structure learning, BCELoss achieves an F1-micro score of $70.74\%$, substantially outperforming the SoftMarginLoss score of 60.77%. This trend is consistent across multiple datasets, highlighting BCELoss as the optimal choice for multi-label adaptation in GCond.
Across all initialization settings, the subgraph sampled by the K-Center strategy receives the best performance in the PPI, Yelp and OGBN-Proteins rating by F1-micro score. However, Random and Herding work quite well among the less dense multi-label datasets like DBLP and PCG. After removing the graph structure, by comparing only the set of nodes of the sampled graphs for label initialization and probability synthetic labels in Table 2, we find that some of the graph structure could have a factor to impact performance like in PPI with 1.00% using K-Center, the result of the subgraph could achieve 47.77% with structure while 36.60% only with nodes. Similarly, K-Center works best among Yelp, OGBN-Proteins, and PCG with node set. Probability synthetic labels capture less of the distribution of the multi-labels in original dataset and only get comparable performance in PPI. After deploying the initialization methods, we also compare these various settings combined with the loss functions discussed. Following the optimization of the condensation process with or without $A^{\prime}$, we report the results of GCond with SoftMarginLoss and BCELoss optimization strategies in Table 3 and Table 4, respectively. Finally, we report the the general initialization methods comparison visually and F1-macro score in appendix A. Interestingly, we find the best performance distribution is consistence even using different loss functions. K-Center initialization with or without graph structure optimization gains the best performance using BCELoss. Noticeably, based on structure learning in Eq.6 the performances gap is quite obvious in PPI and DBLP datasets.
Consistent with Observations 1 and 2, initializing with K-Center sampling and incorporating structure learning $A^{\prime}$ could increase the performance of synthetic multi-label graph. As shown in Observation 3, arcoss a variety of datasets, BCELoss performs better than SoftMarginLoss. Particularly when paired with K-Center and structural learning, BCELoss improves the F1-micro score, suggesting it can offer greater label retention and general adaptability. Therefore, we adopt the K-Center and BCELoss as the initialization and optimization methods with structure information for multi-label scenario.
7 Benchmarking
Table 5: Graph Condensation Methods Performance Comparison with F1-micro and F1-macro. Performance metrics of the model, with the F1-score represented as a decimal value. "-" means the out-of-memory (OOM) errors.
Dataset C-rate Coreset (Random) Coreset (Herding) Coreset (K-Center) SGDD GCDM GCond Whole Dataset
F1-micro F1-macro F1-micro F1-macro F1-micro F1-macro F1-micro F1-macro F1-micro F1-macro F1-micro F1-macro F1-micro F1-macro
PPI 0.50% 41.92 14.69 41.16 12.76 45.65 15.54 44.99 22.58 48.28 23.37 49.63 23.16 51.26 30.06
1.00% 39.07 12.58 41.23 11.66 47.43 17.36 47.02 18.26 47.07 21.76 51.35 23.86
2.00% 41.12 14.12 39.21 10.23 52.34 26.62 46.73 21.24 47.91 21.93 51.21 24.32
PPI-large 0.10% 44.43 14.49 43.17 13.74 41.12 18.11 49.32 18.33 47.13 37.59 51.05 27.51 51.61 28.23
0.20% 45.15 15.80 43.39 15.91 40.88 13.00 44.82 15.78 46.26 22.24 51.99 29.95
0.40% 44.01 15.43 43.61 13.74 41.04 12.89 49.89 29.08 44.08 19.48 52.23 28.59
Yelp 0.01% 33.73 5.55 33.63 7.07 34.20 13.30 - - 28.00 2.78 28.40 7.51 37.97 15.18
0.02% 33.64 5.09 32.94 5.64 35.64 12.30 - - 27.84 3.76 33.94 6.22
0.04% 34.01 5.41 32.87 4.67 36.51 12.51 - - 26.42 2.87 32.77 5.85
0.05% 33.74 5.86 32.83 5.17 36.63 11.30 - - 24.10 2.42 34.11 6.86
0.10% 33.56 5.19 32.83 5.51 36.65 9.79 - - 24.82 2.95 33.70 8.17
0.20% 33.75 5.50 33.20 4.71 37.61 11.48 - - 28.43 3.32 32.41 12.18
DBLP 0.20% 49.28 34.81 41.02 15.41 52.70 49.75 43.17 25.06 44.41 34.84 63.36 56.27 87.55 86.39
0.40% 52.94 44.71 42.59 31.05 55.27 50.89 45.37 43.42 45.88 26.54 68.11 57.35
0.80% 62.08 56.00 41.11 15.63 59.77 56.90 43.48 25.42 46.66 46.20 70.74 68.44
1.60% 68.39 63.75 42.24 25.78 63.87 62.34 42.73 34.22 45.99 26.47 70.49 68.92
3.20% 71.50 68.24 42.70 22.31 67.51 64.54 43.17 25.06 47.81 28.02 70.21 67.78
OGBN-Proteins 0.05% 14.40 1.70 11.29 4.25 28.74 7.83 - - 21.82 7.95 26.38 9.98 10.14 9.16
0.10% 15.63 2.12 15.28 5.21 31.04 10.13 - - 22.18 3.07 29.40 7.47
0.20% 16.90 2.06 8.55 1.64 26.36 4.82 - - 21.38 8.78 29.26 7.71
PCG 2.00% 15.72 5.57 18.98 3.03 27.44 9.09 27.93 15.32 29.48 13.22 18.84 14.21 42.26 31.49
4.00% 15.83 5.20 23.03 13.84 25.13 5.73 28.40 9.00 32.16 8.31 25.58 13.32
8.00% 16.67 3.02 22.48 9.36 18.21 4.59 33.50 8.58 31.88 12.51 26.37 13.90
HumanGo 2.00% 28.88 4.57 19.92 4.39 25.38 6.15 35.58 10.06 36.71 6.53 30.68 11.24 51.67 25.57
4.00% 27.95 7.98 31.70 6.66 31.94 7.24 35.39 9.19 35.79 7.31 34.91 10.74
8.00% 37.57 9.58 32.44 5.50 36.50 7.67 35.25 9.74 36.54 7.35 37.18 12.05
EukaryoteGo 1.00% 31.45 5.57 21.60 4.24 17.79 2.66 29.40 3.25 36.93 4.09 35.24 5.57 45.86 12.27
2.00% 30.72 5.03 26.74 5.16 20.99 3.03 28.69 3.28 36.98 4.09 36.38 7.00
3.00% 30.79 4.18 25.99 3.43 22.08 4.31 36.57 4.75 36.97 4.84 38.90 6.04
In this section, followed the best setting discussed in Section 6, we adapt all the discussed condensation methods into multi-label scenario based on the former settings. The results in Table 5 illustrate the efficacy of different Coreset selection strategies (Random, Herding, K-Center) and graph condensation methods (SGDD, GCDM, GCond) compared to the full dataset. The comparison is performed at different condensation rates (C-rate), providing insights into the impact of these methods on classification performance in multi-label tasks. Specifically, SGDD and GCond both choosing gradient matching for the GNNs trained on original graph $\mathcal{G}$ and synthetic graph $\mathcal{S}$ directly. GCDM uses the embedding matching based on the nodes set belongs to the same class in multi-label.
We have several noteworthy observations concerning the methods, the datasets, the evaluation settings, and the overall results. We highlight the main observations below. Here, the performance metrics of the model, with the F1-score represented as a decimal value.
Observation 1: GCond achieves the best results across datasets. GCond consistently outperforms other methods on most datasets, particularly excelling in large-scale datasets with substantial label complexity. On the PPI dataset, GCond achieves F1-micro and F1-macro scores of 51.35% and 23.86%, respectively, at a 1.00% condensation rate. This robustness extends to PPI-large, where it achieves scores of 52.23% (F1-micro) and 28.59% (F1-macro) at a 0.40% condensation rate. Similarly, on DBLP, GCond demonstrates its superiority with F1-micro and F1-macro scores of 70.74% and 68.44%, respectively, at 0.80% condensation, showcasing its ability to preserve label correlations and structural integrity during condensation. However, GCond’s advantage diminishes in the Yelp dataset at high condensation levels, likely due to the dataset’s unique label structure, which poses challenges for methods relying on label coherence preservation.
Observation 2: K-Center demonstrates stability in extreme condensation setting. The K-Center method shows consistent performance, particularly on large-scale and dense-label datasets, highlighting its strength in preserving diversity at low condensation rates. On the Yelp dataset, K-Center outperforms traditional Coreset methods, achieving the highest F1-micro score (37.61%) at a 0.20% condensation rate. This suggests that its selection strategy effectively retains essential structural and label information, even under high condensation, making it advantageous for large-scale settings where extreme data reduction risks losing critical information.
Observation 3: SGDD’s performance is constrained by scalability issues on larger datasets. SGDD faces out-of-memory (OOM) errors on large datasets like Yelp and OGBN-Proteins, revealing its scalability limitations. However, on smaller datasets, SGDD performs competitively, benefiting from its structure-broadcasting graphon technique, which efficiently captures structural patterns. For example, on datasets such as PCG, HumanGo, and EukaryoteGo, SGDD achieves strong F1-micro and F1-macro scores, demonstrating its effectiveness on smaller scales. Nevertheless, its applicability to larger datasets requires further optimization.
In conclusion, the efficacy of each condensation method varies depending on the characteristics of the dataset. GCond’s superior performance on complex, high-dimensional datasets such as DBLP and PPI demonstrates its strong alignment with datasets that require maintaining label consistency and intricate relationships. In contrast, simpler methods like Random and Herding exhibit relatively stable performance but are outperformed by GCond and K-Center on datasets where preserving structure and label diversity is critical. These findings underscore the importance of selecting methods based on dataset scale, complexity, and label interaction patterns.
8 Conclusion
In this benchmark, we presented a comprehensive framework for multi-label graph condensation, adapting traditional single-label methods to address the complexities inherent in multi-label classification tasks. The results show that the GCond method not only excels in maintaining the structural integrity of the original graph, but also enhances the label correlations, leading to more accurate multi-label predictions.
Appendix A Appendix
A.1 Adapted GCond Algorithm
Input: Training data $\mathcal{G}=(A,X,Y)$
1 Initialize multi-label synthetic graph $\mathcal{S}$ with ${X^{\prime},Y^{\prime}}$
2 for  $k=0$ to $K-1$ do
3       Initialize $\theta_{0}\sim P_{\theta_{0}}$
4       for  $t=0$ to $T-1$ do
5             $D^{\prime}\leftarrow 0$
6             Compute $A^{\prime}=g_{\phi}(X^{\prime})$, then $\mathcal{S}=(A^{\prime},X^{\prime},Y^{\prime})$
7             Compute $\mathcal{L}^{\mathcal{G}}=\mathcal{L}(GNN_{\theta_{t}^{\mathcal{G}}}(A,X),Y)$
8             Compute $\mathcal{L}^{\mathcal{S}}=\mathcal{L}(GNN_{\theta_{t}^{\mathcal{S}}}(A^{\prime},X^{\prime},Y^{\prime}))$
9             $D^{\prime}\leftarrow D^{\prime}+D(\nabla_{\theta_{t}}\mathcal{L}^{\mathcal{G}},\nabla_{\theta_{t}}\mathcal{L}^{\mathcal{S}})$
10
11            if  $t\%(\tau_{1}+\tau_{2})<\tau_{1}$ then
12                   Update $X^{\prime}\leftarrow X^{\prime}-\eta_{1}\nabla_{X^{\prime}}D^{\prime}$
13
14            else
15                   Update $\phi\leftarrow\phi-\eta_{2}\nabla_{\phi}D^{\prime}$
16
17            Update $\theta_{t+1}\leftarrow\text{opt}_{\theta}(\theta_{t},\mathcal{S},\tau_{\theta})$
18
19      Compute new adjacency matrix: $A^{\prime}=g_{\phi}(X^{\prime})$
20       for each edge $(i,j)$ in $A^{\prime}$ do
21             $A^{\prime}_{i,j}\leftarrow A^{\prime}_{i,j}$ if $A^{\prime}_{i,j}>\delta$, otherwise $A^{\prime}_{i,j}\leftarrow 0$
22
23
return $(A^{\prime},X^{\prime},Y^{\prime})$
Algorithm 1 Multi-Label GCond Adaptation
A.2 Visualization of Multi-Label Datasets
In this section, we delve deeper into the characteristics of the condensed graphs by analyzing class distribution and label correlation within the multi-label datasets. This detailed examination is crucial for understanding how effectively the condensation methods preserve the multi-label relationships present in the original datasets.
After obtaining the full benchmark results, we provide a detailed analysis of the condensed graphs by examining the class distribution and label correlation within the synthetic graphs. These statistical properties are crucial for understanding how well the condensed graphs preserve the multi-label relationships present in the original datasets.
  • Label Correlation: To analyze the relationships between labels in the dataset, we define the correlation matrix: Let $G_{\text{labels}}$ be the set of labels from the original dataset $\mathcal{G}$, and $Y$ be the corresponding original label matrix. We denote $M(i,j)$ as the occurrence times of labels $L_{i}$ and $L_{j}$, and $N_{i}$ as the total number of times label $L_{i}$ appears.
    The conditional probability $P(i,j)$ is defined as:
    $P(i,j)=\frac{M(i,j)}{N_{i}}$ (24)
    This conditional probability matrix $P$ models the edges as a co-occurrence matrix.
    Next, we define the diagonal matrix $D$ as follows:
    $D=\text{Diag}(P)$ (25)
    Finally, the Laplacian matrix $L_{o}$ is computed as:
    $L_{o}=D-P$ (26)
    By computing the label correlation matrix, we investigate whether the co-occurrence patterns between labels in the original graph are retained in the condensed graph.
  • Class Distribution: We analyze how evenly or unevenly the labels are distributed across the classes in different datasets by index, which helps assess the quality of the condensation process.
Following the above definition, we report the visualizations of different datasets as follows. The label correlation and class distribution are shown in Figures 2 and 3, respectively. We find that the more complex the labels, the more the condensation methods would rely on structure and become more random. For future work, more suitable methods need to be fit to multi-label scenarios.
(a) PPI
(b) PPI-Large
(c) Yelp
(d) OGNB-Proteins
(e) DBLP
(f) PCG
(g) HumanGO
(h) EukaryoteGo
Figure 2: Multi-label Correlation Visualization
(a) PPI
(b) PPI-Large
(c) Yelp
(d) OGBN-Proteins
(e) DBLP
(f) PCG
(g) HumanGO
(h) EukaryoteGo
Figure 3: Multi-label Class Distribution Visualization
A.3 F1-Macro Results of Adaptation
Figures 4, 5 and 6 show the visualizations of different adapation strategies measured by F1-micro score. To further investigate the effectiveness of the methods, we also report the F1-macro score in Tables 6 , 8 and 8.
Figure 4: Different Initialization Methods Performance. Performance metrics of the model, with the F1-score represented as a decimal value.
Figure 5: Performance Optimized by SoftMargin Loss. Performance metrics of the model, with the F1-score represented as a decimal value.
Figure 6: Performance Optimized by BCE Loss. Performance metrics of the model, with the F1-score represented as a decimal value.
Datasets C-rate Random Herding K-Center Probability Whole Dataset
Subgraph Nodes Subgraph Nodes Subgraph Nodes
PPI 1.00% 12.58 10.67 11.66 17.98 17.36 18.37 12.55 30.06
Yelp 0.02% 5.09 3.76 5.64 3.66 12.30 8.78 2.87 15.18
DBLP 0.80% 56.00 51.43 15.63 15.41 56.90 52.74 15.68 86.39
OGBN-Pro 0.1% 2.12 7.12 5.21 7.17 10.13 6.42 6.36 9.16
PCG 4% 5.20 9.69 13.84 6.45 5.73 7.10 11.35 31.49
Table 6: F1 Macro Score(%) of Coreset Method with Different Initializetion Strategies
Datasets C-rate Random Herding K-Center Probability Whole Dataset
Without $A^{\prime}$ With $A^{\prime}$ Without $A^{\prime}$ With $A^{\prime}$ Without $A^{\prime}$ With $A^{\prime}$ Without $A^{\prime}$ With $A^{\prime}$
PPI 1.00% 22.60 28.33 30.37 18.32 11.62 23.12 16.66 12.61 30.06
Yelp 0.02% 3.72 6.98 3.57 3.99 6.37 9.02 2.13 1.86 15.18
DBLP 1% 35.81 37.84 27.08 27.11 32.21 49.36 42.33 38.08 86.39
OGBN-Pro 0.10% 4.36 5.47 5.37 7.18 5.26 5.32 3.21 4.76 9.16
PCG 4% 12.42 13.86 10.09 15.50 7.34 9.47 13.29 6.50 31.49
Table 7: F1-Macro Score (%) of GCond Method with Random/Herding/K-Center/Probability Distribution Initialization with/without Learning from Graph Structure for SoftMarginLoss
Datasets C-rate Random Herding K-Center Probability Whole Dataset
Without $A^{\prime}$ With $A^{\prime}$ Without $A^{\prime}$ With $A^{\prime}$ Without $A^{\prime}$ With $A^{\prime}$ Without $A^{\prime}$ With $A^{\prime}$
PPI 1.00% 22.82 23.86 21.72 21.35 20.69 21.47 17.68 20.62 30.06
Yelp 0.02% 3.57 8.02 4.24 4.13 6.13 6.22 4.00 3.07 15.18
DBLP 1% 34.89 51.44 26.94 28.28 34.87 68.44 38.36 39.87 86.39
OGBN-Pro 0.10% 6.79 6.19 5.87 7.30 4.98 7.47 2.89 4.99 9.16
PCG 4% 7.83 7.29 15.17 14.27 8.02 13.32 13.13 11.30 31.49
Table 8: F1-Macro Score (%) of GCond Method with Random/Herding/K-Center/Probability Distribution Initialization with/without Learning from Adjacent Nodes for BCELoss
A.4 Class-Weighted Optimization
Let $N_{\max}=\max_{c\in{0,\cdots,C-1}}N_{c}$ is the maximum number of samples across all classes. For each class $c$ the class-wise coefficient is defined as:
$\alpha_{c}=\frac{N_{c}}{N_{\max}}$ (27)
The total loss for all the $N$ nodes would be the sum over all the classes and their respective nodes $N_{c}$:
$\mathcal{L}_{single-label}=\sum_{c=0}^{C-1}\alpha_{c}\sum_{j\in N_{c}}\ell_{CE}(z_{j},y_{j})$ (28)
This is the general equation for different methods, as the condensation part is using the $\mathcal{M}(\cdot)$ as the matching loss, for example, GCond would compute the gradient for each class with class-wise coefficient and sum them up to get the final loss.
Furthermore, for each task we are working on the binary classification problem. To better capture the contribution of each class in multi-label, we introduce class weights $\omega\in\mathbb{R}^{K}$ for each task $k$:
$\ell_{\text{BCE}}(z_{j,k},y_{j,k})=-\omega_{k}(y_{j,k}\log(\sigma(z_{j,k}))+(1-y_{j,k})\log(\sigma(z_{j,k}))),$ (29)
where $\omega_{k}$ is the positive class weight, scaling the loss for each positive class $y_{i,k}=1$. Therefore, the final loss of multi-label classification is:
$\mathcal{L}_{multi-label}=-\frac{1}{N}\sum_{j=0}^{N-1}\sum_{k=0}^{K-1}\omega_{k}(y_{j,k}\log(\sigma(z_{j,k}))$ (30)
$+(1-y_{j,k})\log(1-\sigma(z_{j,k})))$
This is helpful when there is an imbalance between the number of positive and negative samples. The F1-micro results are shown in Table 9 and Figure 7. F1-macro results are shown in Table 10.
Figure 7: Performance Optimized by BCE Loss with Class Weight $\omega$
Datasets C-rate Random Herding K-Center Probability Whole Dataset
Without $A^{\prime}$ With $A^{\prime}$ Without $A^{\prime}$ With $A^{\prime}$ Without $A^{\prime}$ With $A^{\prime}$ Without $A^{\prime}$ With $A^{\prime}$
PPI 1.00% 49.30 51.28 46.25 46.65 48.10 48.15 46.32 43.64 51.26
Yelp 0.02% 31.06 30.12 32.42 32.58 34.26 33.75 22.10 25.82 37.97
DBLP 0.80% 52.47 65.27 46.19 48.79 52.58 60.09 50.43 52.69 87.55
OGBN-Proteins 0.10% 26.51 23.25 22.25 26.36 24.81 28.24 26.47 30.13 30.59
PCG 4% 23.78 25.72 18.36 23.53 21.15 24.10 24.68 30.13 42.26
Table 9: F1-Micro Score (%) of GCond Method with Random/Herding/K-Center/Probability Distribution Initialization with/without Learning from Adjacent Nodes for BCE Loss and Balanced Coefficient
Datasets C-rate Random Herding K-Center Probability Whole Dataset
Without $A^{\prime}$ With $A^{\prime}$ Without $A^{\prime}$ With $A^{\prime}$ Without $A^{\prime}$ With $A^{\prime}$ Without $A^{\prime}$ With $A^{\prime}$
PPI 1.00% 21.75 25.65 15.55 27.21 27.09 31.97 14.31 21.43 30.06
Yelp 0.02% 3.91 6.49 4.06 5.16 11.72 7.78 1.86 2.42 15.18
DBLP 0.80% 32.79 61.83 26.44 29.25 32.64 56.29 37.78 42.11 86.39
OGBN-Pro 0.10% 4.60 6.05 3.54 4.69 6.42 7.89 4.38 8.56 9.16
PCG 4% 13.26 11.93 12.24 15.75 5.70 9.88 11.60 8.56 31.49
Table 10: F1-Macro Score (%) of GCond Method with Random/Herding/K-Center/Probability Distribution Initialization with/without Learning from Adjacent Nodes for BCE Loss and Balanced Coefficient